OpenStudy (anonymous):

The approximation of the natural logarithm of 2: ln 2 ≈ 0.693 is commonly used by applied scientists, biologists, chemists, and computer scientists. For example, chemists use it to compute the half-life of decaying substances. Based on this approximation and the power rule for logarithmic expressions, how could you approximate ln 8, without a calculator? Explain.

OpenStudy (lgbasallote):

hint: \[\ln 8 \implies \ln (2^3) \implies 3\ln 2\] does that help?

OpenStudy (anonymous):

This helps a lot. Thanks.

OpenStudy (lgbasallote):

welcome